7, 21, 8, 72, 9, ? What completes the sequence 81 243 123 136

Answer by Edwin McCravy(20056)

7, 21, 8, 72, 9, ?
1st term = 7
2nd term = 7×31 = 7×3 = 21
3rd term = 7+1 = 8
4th term = 8×32 = 8×9 = 72
5th term = 8+1 = 9
6th term = 9×33 = 9×27 = 243
7th term = 9+1 = 10
Edwin